OpenStudy (anonymous):

Could you check my answer? y = x^2 + 3x + 2 List all the factors of 2: 1 * 2 = 2 -1 * -2 = 2 Find the product of 2 and sum of 3. 1 * 2 = 2 and 1 + 2 = 3 Place the numbers into a binomial: (x + 2)(x + 1) Is that right? It just seems too simple to be the right answer.

jimthompson5910 (jim_thompson5910):

you are correct, x^2 + 3x + 2 factors to (x+2)(x+1)

OpenStudy (anonymous):

to check yourself, all you gotta do is multiply the factors back together.

jimthompson5910 (jim_thompson5910):

Check: (x+2)(x+1) = x(x+1) + 2(x+1) (x+2)(x+1) = x*x+x*1 + 2*x+2*1 (x+2)(x+1) = x^2+x + 2x + 2 (x+2)(x+1) = x^2 + 3x + 2

OpenStudy (anonymous):

your right @haileemackk

OpenStudy (anonymous):

Okay, thanks! I wasn't completely sure.

OpenStudy (anonymous):

:) Your welcome